OpenStudy (anonymous):

one roots of x^2 - lx + m =0 is the square of the other . Then m^2 + 3lm + m=?

OpenStudy (experimentx):

use the same technique as before.

OpenStudy (experimentx):

rest is just algebraic manipulation

OpenStudy (anonymous):

let y be the root eq1 and y^2 be the root of 2nd eq

OpenStudy (experimentx):

let y and y^2 be root of the given eqn.

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

sum = y + y^2 = L product = y^3 =M

OpenStudy (experimentx):

yep .. manipulate around those things.

OpenStudy (anonymous):

i am not getting ...

OpenStudy (anonymous):

i got such eq y^6 + 3y^5 + 3Y^4 + y^3

OpenStudy (experimentx):

do you know the answer?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

it should be l^3

OpenStudy (anonymous):

L^3

OpenStudy (experimentx):

(y^2+y)^3 = ?

OpenStudy (anonymous):

l^3 thxxx

OpenStudy (experimentx):

i told you so ... if you identify the problem correctly ... the problem is 50% solved if you know the answer \75% solved.